# Standing on the Shoulders of Giants: Reprogramming Visual-Language Model for General Deepfake Detection

### v1

@@ Introduction
Cross-Datasets Evaluations
The cross-datasets evaluation is still a challenging task because the unknown domain gap between the training and testing datasets can be caused by different source data, forgery methods, and/or post-processing. In this part, we evaluate the generalization performances in a cross-dataset setting. Specifically, our models were trained on the FF++ (only containing DF, F2F, FS, and NT subsets for fair comparisons) and tested on other datasets. Our method is compared with several state-of-the-art (SOTA) methods published in the past three years, including: UiA-ViT(Zhuang et al. 2022), DCL(Sun et al. 2022), CFM (Luo et al. 2023), AUNet(Bai et al. 2023), SLADD(Chen et al. 2022), SBI(Shiohara and Yamasaki 2022), FoCus(Tian et al. 2024), UCF(Yan et al. 2023), TALL++(Xu et al. 2024), TALL(Xu et al. 2023), (Ba et al. 2024), SeeABLE(Larue et al. 2023), LSDA(Yan et al. 2024), LAA-Net(Nguyen et al. 2024), VLFFD(Sun et al. 2023), IID(Huang et al. 2023), SA3WT(Li et al. 2024), and Bi-LIG(Jiang et al. 2024). The experimental results in terms of frame-level and video-level AUC are shown in Table 2. Aside from its moderate performance on the CDF dataset, our method outperforms most competitors on the Wild, DFDC, and DFDCP datasets, which are closer to real-world scenarios. We also report the results trained on FF++/DF, and found our method still performs better than all the competitors on the Wild, DFDC, and DFDCP datasets. It is important to highlight that all competitors re-train backbone networks (e.g., ResNet18 and ViT), which consist of at least 10M trainable parameters. In contrast, our method utilizes a mere 0.078M parameters, illustrating that it can achieve superior general performance with fewer learnable parameters.
Cross-Manipulations Evaluations
Existing face forgery detectors often struggle to handle emerging manipulation techniques. In this part, we conduct cross-manipulation experiments involving five forgery techniques, i.e., Deepfakes (DF), Face2Face (F2F), FaceSwap (FS), and NeuralTextures (NT), FaceShifterFSh. We examine models trained on one manipulation type and tested across the other three. As shown in Table 3, we can observe that our method can improve cross-manipulation performances. Note that for the face reenactment (i.e., F2F and NT), face identity is not changed, thus the performance is a little worse than SOTA. Therefore, Face ID plays a limited guiding role, which leads to limited improvement of our method. These results highlight the effectiveness of our method in combating emerging unseen forgery methods.
Evaluations of RepDFD
In this part, we perform several evaluations to explore the effectiveness of ReDFD. The main results of the following experiments are cross-datasets performances trained on FF++/DF.
@@
In this part, we evaluate the effectiveness of our reprogramming paradigm compared with other tuning paradigms. Specifically, for the fine-tuning of the image feature encoder $E_{I}$, two established methods were assessed: Full Fine-Tuning (FFT), entailing the adjustment of all parameters within $E_{I}$, and Linear Probing (LP), which incorporates a learnable linear layer while maintaining $E_{I}$ as frozen. In addition, we reference a very recent work MoE-FFD(Kong et al. 2024), which presents a deepfake detection method jointly utilizing the LoRA and Adapter paradigms to tune a frozen $E_{I}$.
Beyond the image feature extractor-related tuning paradigms, our study also incorporated a text feature extractor-related tuning paradigm, CoOp(Zhou et al. 2022), into the comparison. This method maintains the CLIP frozen while introducing learnable text prompts to adapt CLIP for specific target tasks. As shown in Table 4, our approach outperforms other methods in most scenarios. LP and CoOp inadequately facilitate the transfer of the base CLIP model to the deepfake detection task, resulting in suboptimal generalization performance. We think that these two paradigms exhibit a shared issue; they influence the utilization of image features, rather than the extraction of these features. Compared with them, FFT, MoE-FFD, and ours can directly impact image feature extraction, consequently improving general performance across multiple datasets. Furthermore, our method achieves superior generalization using significantly fewer parameters. We speculate that utilizing a limited number of parameters to fit deepfake-related knowledge potentially preserves the generalization capabilities of the base model. Notably, the performance gain is substantial when using a larger CLIP model with our method. Conversely, the performance of MoE-FFD significantly declines when utilizing a larger model, suggesting possible overfitting. It reveals the potential of our approach to be effectively scaled to larger models.
@@
In this part, we investigate the effects of various text prompt configurations on RepDFD, encompassing fixed text prompts, randomly initialized text prompts, and our adaptive face-related text prompts (termed dynamic text prompts). It can be concluded from the experiment that the dynamic text prompts are more effective than the fixed those. We consider all groups of the real/fake text templates in Table 1, i.e., $\{T_{0},T_{1}\},\{T_{2},T_{1}\},\{T_{2},T_{3}\}$. Furthermore, we consider a special text prompt setting, named ‘Rand Text’, which contains two completely random initialized and frozen text embedding as long as the ‘Fixed Text’ ($\{T_{0},T_{1}\}$). As shown in Table 5, the best performance occured when the face embeddings are only utilized in the text prompt for the fake class, corresponding to the dynamic text prompts setting $\{T_{0},T_{3}\}$. We observed that the ‘Fixed Text’ ($\{T_{0},T_{1}\}$) demonstrated limited effectiveness, performing similarly to ‘Rand Text’, which lacks substantive semantic content. This observation implies that semantic content in language may not significantly influence deepfake detection tasks, as the focus primarily lies on trace analysis. In contrast, integrating face embeddings into text prompts to create dynamic text prompts, corresponding to $\{T_{2},T_{3}\},\{T_{2},T_{1}\},\{T_{0},T_{3}\}$, boosted model performance. We posit that the operation introduces fine-grained and face-ralated visual information into text prompts, thereby incorporating details that are challenging to describe linguistically. Therefore, the dynamic text prompts can not only enrich the supervision information during training, but also provide sample-level adaptive information to support classification.
@@
Impact for different face embeddings
To further verify the universality of Face2Text, we investigate the impact of different face embeddings on Face2Text prompts. We perform an ablation study comparing the results obtained by ArcFace(Deng et al. 2019), BlendFace(Shiohara, Yang, and Taketomi 2023) and Transface(Dan et al. 2023) and non-face (i.e., using the text prompt group $\{T_{0},T_{1}\}$). As shown in Figure 5, our method demonstrates good performance across vairous face encoders $F_{id}$.
Discussion
In this subsection, we present several in-depth analyses below, aiming to better understand the effectiveness of our method.
@@ Discussion
In Table 5, it is noteworthy that optimal performance is achieved when the $\{T_{0},T_{3}\}$ configuration is selected, which solely incorporates face embedding into the text for the fake label. To investigate the reason, as shown in Table 6, we calculated the cosine similarity between the image features extracted by the initialized visual prompt $\boldsymbol{\delta}$ and the text prompts, both with and without face embedding. Our findings indicate that cosine similarities tend to be higher for the real label ($T_{0}$) when using the text prompt without face embeddings, and for the fake label ($T_{3}$) when using the text prompt with face embeddings. We guess that text prompts with higher cosine similarities offer a better initialization for the CLIP model, enabling more effective fine-tuning of target models. This observation may benefit future methodological design.
The visualization of feature distribution
@@
In this experiment, we provide the visualization of feature distributions as shown in Figure 6. The influence of the visual prompt $\boldsymbol{\delta}$ is limited exclusively to the input images and not the models, ensuring that the image features can occupy in a consistent feature latent space. Initial observations suggest that there are significant differences in the orginal feature distributions between the CDF and FF datasets, which indicates the intrinsic domain discrimination capability of the unmodified CLIP. Moreover, before integrating visual prompts, the original CLIP model failed to differentiate between real and fake images. In contrast, within the FF dataset, real and fake images can be efficiently discriminated by equipping visual prompts, suggesting the effectiveness of our method. Furthermore, although the CDF dataset was not used during training, visual prompts trained on the FF dataset effectively endowed the model with the capability to detect deepfakes in the CDF dataset. We speculate that the visual prompts $\boldsymbol{\delta}$, consisting of a limited number of adjustable parameters, potentially tend to exploit the inherent capabilities of frozen models instead of introducing additional deepfake-information directly, which maybe protect the generalization ability of pretrained models so that improve the peformance. We think the speculation is necessary to further explore in future works.
@@
Figure 6: The t-SNE visualization of image features. ’CLIP’ denotes features extracted by the original CLIP feature encoder, while ’CLIP+VP’ refers to features extracted by the visual prompt.
Conclusions and Discussions
@@
In this paper, we propose RepDFD, a general yet parameter-efficient method for detecting face forgeries by reprogramming a well-trained CLIP model. Specifically, we employ an Input Transformation technique to merge the image with learnable perturbations before feeding it into the CLIP image encoder. Additionally, we introduce Face2Text Prompts to incorporate facial identity information into the text prompts, which are then fed into the CLIP text encoder to guide the optimization of perturbations. The only required operation for evaluation is to apply these task-specific perturbations to all test images and then feed them into the CLIP model. Comprehensive experiments show that: (1) the cross-dataset and cross-manipulation performances of deepfake detection can be significantly and consistently improved using a pre-trained CLIP model with RepDFD; (2) our superior performances are at less cost of trainable parameters, making it a promising approach for real-world applications.

### v4

@@ Introduction
Cross-Dataset Evaluations
The cross-datasets evaluation is still a challenging task because the unknown domain gap between the training and testing datasets can be caused by different source data, forgery methods, and/or post-processing. In this part, we evaluate the generalization performances in a cross-dataset setting, in which detection models were trained on the FF++ (only containing DF, F2F, FS, and NT subsets for fair comparisons) and tested on other datasets. Our method is compared with several state-of-the-art (SOTA) methods proposed in the past three years, including: UiA-ViT (Zhuang et al. 2022), DCL (Sun et al. 2022), CFM (Luo et al. 2023), AUNet (Bai et al. 2023), SLADD (Chen et al. 2022), SBI (Shiohara and Yamasaki 2022), FoCus (Tian et al. 2024), UCF (Yan et al. 2023a), TALL++ (Xu et al. 2024), TALL (Xu et al. 2023), (Ba et al. 2024), SeeABLE (Larue et al. 2023), LSDA (Yan et al. 2024a), LAA-Net (Nguyen et al. 2024), VLFFD (Sun et al. 2023), IID (Huang et al. 2023), SA3WT (Li et al. 2024), and Bi-LIG (Jiang et al. 2024). The experimental results in terms of frame-level and video-level AUC are shown in Tab. 2. Aside from its moderate performance on the CDF dataset, our method outperforms most competitors on the Wild, DFDC, and DFDCP datasets. We also report the results trained on FF++/DF, and found our method still performs better than all the competitors on the Wild, DFDC, and DFDCP datasets. It is important to highlight that all competitors re-train backbone networks (e.g., ResNet18 and ViT), which consist of at least 10M trainable parameters. In contrast, our method utilizes a mere 0.078M parameters, illustrating its superior general performance with fewer learnable parameters.
Cross-Manipulation Evaluations
Existing face forgery detectors often struggle to handle emerging manipulation techniques. In this part, we conduct cross-manipulation experiments involving three forgery techniques, i.e., Deepfakes (DF), FaceSwap (FS), and FaceShifter (FSh). We examine models trained on one manipulation type and tested across the other three. As shown in Tab. 3, it can be observed that our method can improve cross-manipulation performances. These results highlight the effectiveness of our method in combating emerging unseen forgery methods.
Ablation Studies
In this part, we perform several evaluations to explore the effectiveness of ReDFD. The main results of these experiments are cross-dataset performances trained on FF++/DF.
@@
In this part, we evaluate the effectiveness of our reprogramming paradigm compared with other tuning paradigms. Specifically, for the fine-tuning of the image feature encoder $E_{I}$, two established methods were assessed: Full Fine-Tuning (FFT), entailing the adjustment of all parameters within $E_{I}$, and Linear Probing (LP), which incorporated a learnable linear layer while maintaining $E_{I}$ as frozen. In addition, we reference a very recent work MoE-FFD (Kong et al. 2024), which presents a deepfake detection method jointly utilizing the LoRA and Adapter paradigms to tune a frozen $E_{I}$.
Beyond these three image feature extractor-related tuning paradigms, our study also incorporated a text feature extractor-related tuning paradigm, CoOp (Zhou et al. 2022), into the comparison. It maintains the CLIP frozen while introducing learnable text prompts to adapt CLIP for target tasks. As shown in Tab. 4, our approach outperforms other methods in most scenarios. LP and CoOp inadequately facilitate the transfer of the base CLIP model to the deepfake detection task, resulting in suboptimal generalization performance. We consider that these two paradigms share a common issue: they affect the utilization of image features rather than their extraction. Compared with them, FFT, MoE-FFD, and ours can directly impact image feature extraction, consequently improving general performance across multiple datasets. Furthermore, our method achieves superior generalization using significantly fewer parameters. Although the number of trainable parameters does not increase, the performance gains when using a larger CLIP model with our method. Conversely, the performance of MoE-FFD significantly declines on a larger model due to more trainable parameters, suggesting possible overfitting. We speculate that utilizing a limited number of parameters to fit deepfake-related knowledge potentially preserves the generalization capabilities of the base model. It reveals the potential of our approach to be effectively scaled to larger models.
@@
In this part, we investigate the effects of various text prompt configurations on RepDFD, including fixed text prompts, randomly initialized text prompts, and our adaptative face-related text prompts (termed dynamic text prompts). It can be concluded from the experiment that the dynamic text prompts are more effective than the fixed those. We consider all groups of the real/fake text templates in Tab. 1, i.e., $\{T_{0},T_{1}\},\{T_{2},T_{1}\},\{T_{2},T_{3}\}$. Besides, we consider a special text prompt setting, named ‘Rand Text’, which contains two completely random initialized and frozen text embedding as long as the ‘Fixed Text’ ($\{T_{0},T_{1}\}$). As shown in Tab. 5, the best performance occurred when the face embeddings were only utilized in the text prompt for the fake class, corresponding to the dynamic text prompts setting $\{T_{0},T_{3}\}$. We observed that the ‘Fixed Text’ ($\{T_{0},T_{1}\}$) demonstrated limited effectiveness, performing similarly to ‘Rand Text’, which lacks substantive semantic content. This observation implies that semantic content in language may not significantly influence deepfake detection tasks, as the primary focus is on trace analysis. In contrast, integrating face embeddings into text prompts to create dynamic text prompts, corresponding to $\{T_{2},T_{3}\},\{T_{2},T_{1}\},\{T_{0},T_{3}\}$, boosted model performance. It may introduce fine-grained and face-related visual information into text prompts, thereby supplementing details that are challenging to describe linguistically. Thus, the dynamic text prompts can not only enrich the supervision information during training, but also provide sample-level adaptative information to support classification.
@@
Impact of different face embeddings
To further verify the universality of Face2Text, we investigate the impact of different face embeddings on Face2Text prompts. We compare the results obtained by ArcFace (Deng et al. 2019), BlendFace (Shiohara, Yang, and Taketomi 2023) and Transface (Dan et al. 2023) and fixed text (i.e., using the text prompt group $\{T_{0},T_{1}\}$). As shown in Fig. 5, our method demonstrates good performance across various face encoders $E_{F}$.
@@
Discussion
In this subsection, we present several in-depth analyses below, aiming to explore the effectiveness of our method.
@@ Discussion
In Tab. 5, it is noteworthy that optimal performance is achieved when the $\{T_{0},T_{3}\}$ configuration is selected, which solely incorporates face embedding into the text for the fake label. To investigate the reason, as shown in Tab. 6, we calculated the cosine similarity between the image features extracted by the initialized visual prompt $\boldsymbol{\delta}$ and the text prompts, both with and without face embedding. Our findings indicate that cosine similarities tend to be higher for the real label ($T_{0}$) when using the text prompt without face embeddings, and for the fake label ($T_{3}$) when using the text prompt with face embeddings. We speculate that text prompts with higher cosine similarities offer a better initialization for the CLIP model, enabling more effective fine-tuning of target models. This observation may benefit future methodological design.
The visualization of visual feature distribution
@@
In this experiment, we provide the t-SNE visualization of visual feature distributions as shown in Fig. 6. The influence of the visual prompt $\boldsymbol{\delta}$ is limited exclusively to the input images and not the models, ensuring that the image features can occupy a consistent feature latent space. Initial observations suggest that there are significant differences in the original feature distributions between the CDF and FF datasets, which indicates the intrinsic domain discrimination capability of the unmodified CLIP. Moreover, without visual prompts, the original CLIP model failed to identify real and fake images. In contrast, within the FF dataset, real and fake images can be efficiently discriminated by equipping visual prompts, suggesting the effectiveness of our method. Furthermore, although the CDF dataset was not used during training, visual prompts trained on the FF dataset effectively endowed the model with the capability to detect deepfakes in the CDF dataset. We speculate that the visual prompt $\boldsymbol{\delta}$, consisting of a limited number of adjustable parameters, potentially tends to exploit the inherent capabilities of frozen models instead of introducing additional deepfake information directly, which may protect the generalization ability of pre-trained models and improve models’ performance. Except for the observation of visual features, we also provide the visualization of text features and face embeddings in our appendix (Lin et al. 2024a), which reveals a mid-domain between different datasets in common. These observations are necessary to be further explored in future works.
@@
Figure 6: The t-SNE visualization of image features. ‘CLIP’ denotes features extracted by the original CLIP, while ‘CLIP+VP’ refers to features extracted by the visual prompt.
Conclusion
@@
In this paper, we have proposed RepDFD, a general but parameter-efficient method for detecting face forgeries by reprogramming a well-trained CLIP model. Specifically, we employ the Input Transformation to merge the image with learnable perturbations before feeding it into the CLIP image encoder. Moreover, we introduce the Face2Text Prompts to asymmetrically incorporate facial embedding information into the text prompts for real and fake categories, which are then fed into the CLIP text encoder to guide the optimization of perturbations. RepDFD has effectively employed a CLIP to detect deepfakes by processing only the input images and texts, excluding the internal model parameters. Comprehensive experiments have demonstrated that our superior performance can be achieved with fewer trainable parameters.
Appendix
We elaborate on more details and results of our work in this supplementary material.
More Details of Settings
More Implementation Details
In the pre-processing stage, for every video frame in datasets, we employed MTCNN(Zhang et al. 2016) to detect and crop the facial regions, enlarged by a factor of 1.3, and subsequently resized them to 224$\times$224.
All of the code and pre-trained models of CLIP are stemmed from the official repository OpenCLIP(Ilharco et al. 2021). For extracting face embeddings, we employed Transface(Dan et al. 2023) as the default face encoder in our method, using ViT-L version pre-trained on Glint360K(An et al. 2022). It should be noticed that the dimension of face embeddings is $512$. For CLIP-ViT-B, the face embedding can directly be integrated into text embeddings, due to the naturally alignment of dimension. For CLIP-ViT-L, the text embedding dimension is $768$, which presents a feature integration mismatch issue. To address the issue, a random projection layer was implemented to project the face embeddings into the target dimension. The random projection layer was initialized using the function torch.nn.init.normal(mean=0, std=1/768).
During training, the training batch size was set to $32$ and our method did not utilize any data augmentations. Notably, to enable Mixed Precision Training, our models were trained based on the Python library torch.cuda.amp.
More Details of Datasets
We conduct evaluations on widely-used datasets and follow previous settings used in their corresponding datasets and compare with other methods respectively. More details on these datasets are described below.
  • CelebDF (CDF) (Li et al. 2020c) contains 590 real videos of 59 celebrities and corresponding 5639 high-quality fake videos generated by an improved forgery method. We use the stand test set consisting of 518 videos for our experiments.
  • DeepFake Detection Challenge Preview (DFDCP) (Dolhansky et al. 2019) is generated by two kinds of synthesis methods on 1131 original videos. We use all 5250 videos for our experiments.
  • DeepFake Detection Challenge (DFDC) (Dolhansky et al. 2020) is widely acknowledged as the most challenging dataset due to containing many manipulation methods and perturbation noises. We use the public test set consisting of 5000 videos for our experiments.
  • WildDeepfake (Wild) (Zi et al. 2020) contains 3805 real face sequences and 3509 fake face sequences collected from Internet. Thus, it has a variety of synthesis methods and backgrounds, as well as character identities. We use the stand test set consisting of 806 sequences for our experiments.
More Experiments
Cross-Dataset Evaluations
To comprehensively show the performance comparisions, we further supplement the results with the equal error rate (EER) metrics, which represents the point on the Receiver Operating Characteristic (ROC) curve where the false positive rate equals the false negative rate, providing a balanced measure of classification performance. As shown in Tab. 7, the results also show promissing performance like the results with AUC metrics. Our method’s EER performance shows a consistent improvement compared to the AUC metric, highlighting the effectiveness of our approach.
Imapct of the size of the visual prompt
In our method, we incorporate visual prompts with input images processed through Input Transformation. Thus, the border width $p$ of visual prompts impacts both performance and the number of learnable parameters. Herein, we conducted an ablation study to explore the imapct of $p$ was varied among the values $\{12,23,34,45,56,67,78\}$. This variation corresponds to resizing the input images to $\{90\%,80\%,70\%,60\%,50\%,40\%,30\%\}$ of their original size, and then pad it to original size by merging the visual prompt. Tab. 8 exhibits the impact of varying $p$. Intuitively, the learnable parameters of the visual prompt increase with an increase in $p$. However, it is worth noting that as $p$ increases, the average generalization performances of the model initially improves but then drops significantly. We speculate that such decline can be attributed to information loss caused by excessive scaling down of the images, suggesting a necessary trade-off between learnable parameters and the size of the input images. Therefore, we set $p=34$ in our experiments, corresponding to a resized image that is $70\%$ of its original size.
Frame-level Video-level
Method CDF Wild DFDCP DFDC Method CDF Wild DFDCP DFDC
UIA-ViT (ECCV 2022) - - - - DCL (AAAI 2022) 19.12 31.44 29.55 30.94
CFM (TIFS 2024) 24.74 30.79 - 31.67 AUNet (CVPR 2023) - - - -
SLADD (CVPR 2022) - - - - SBI (CVPR 2022) 19.41 37.63 25.00 35.27
FoCus (TIFS 2024) - - - - TALL (ICCV 2023) - - - -
UCF (ICCV 2023) - - - - TALL++ (IJCV 2024) - - - -
Ba et al. (AAAI 2024) - - - - SeeABLE (ICCV 2023) - - - -
LSDA (CVPR 2024) - - - - LAA-Net (CVPR 2024) - - - -
VLFFD (arXiv 2023) 22.73 24.40 23.43 - IID (CVPR 2023) - - - -
SA3WT (IJCV 2024) - - - - Bi-LIG (TIFS 2024) 7.30 - 17.03 25.07
Ours (DF) 28.99 21.04 23.29 34.27 Ours (DF) 20.22 20.45 17.77 30.00
Ours (FF++) 27.44 21.60 18.03 30.32 Ours (FF++) 17.98 20.71 13.00 28.32
Table 7: EER (%) of cross-datasets evaluations. The results of other SOTA methods are directly cited from their corresponding original paper. The best results are highlighted.
The visualization of textual feature distribution
In this experiment, we provide the t-SNE visualization of textual feature distributions. In RepDFD, textual features are generated using text templates and facial embeddings. As illustrated in Fig. 7, the distributions of facial embeddings significantly do not overlap between the FF and CDF datasets due to the differences across domains. However, in Fig. 8, we observe that the distributions of textual features (corresponding to $T_{3}$) for FF and CDF are more closely aligned, forming a common mid-domain. Therefore, we speculate that the textual encoder $E_{T}$ can squeeze different domains into a common mid-domain, which can effectively enhance the learning of our visual prompt $\delta$ by reducing domain discrepancies.
Figure 7: The t-SNE visualization of facial features in FF and CDF datasets.
Figure 8: The t-SNE visualization of textual features in FF and CDF datasets.
$p$Para CDF Wild DFDCP Avg
12 0.031 M 85.28 85.25 89.82 86.78
23 0.055 M 85.55 87.01 90.88 87.81
34 0.078 M 88.41 87.73 90.68 88.94
45 0.097 M 80.64 82.68 90.37 84.56
56 0.113 M 82.17 85.99 91.40 86.49
67 0.126 M 78.59 80.05 88.65 82.43
78 0.137 M 74.18 70.10 86.90 77.06
Table 8: The generalization performance involves different $p$ of the visual prompt. All models were trained on FF++ (DF).